# arXiv:2603.00824v1[cs.LG] 28 Feb 2026

### v1

@@ Abstract
We prove and empirically instantiate four results on a frozen open LLM (Llama 3.2 3B Instruct) using standard text and code corpora (WikiText-103 raw, a C4-derived English web-text subset, and the-stack-smol) Meta (2024); Merity et al. (2016); Raffel et al. (2020); Kocetkov et al. (2022). (A) A constructive spanning-tree gauge theorem shows that, after gauge fixing on a spanning tree, each chord residual equals the holonomy of its fundamental cycle, making holonomy computable and gauge-invariant. (B) Shearing lower-bounds a data-dependent transfer mismatch energy, turning $D_{\mathrm{shear}}$ into an unavoidable failure bound. (C) A non-vacuous certified jamming/interference bound holds on consequential atom subsets with high coverage and zero violations across random seeds and hyperparameters. (D) Bootstrap and sample-size experiments establish estimator stability for $D_{\mathrm{shear}}$ and $D_{\mathrm{hol}}$, with improved concentration on persistent (well-conditioned) subsystems predicted by polar-factor perturbation theory Higham (1986); Schönemann (1966). Our framework supplies reproducible evaluation primitives for when and why global interpretability breaks, separating failures due to local overload (jamming), cross-chart mismatch (shearing), and intrinsically path-dependent transport (holonomy).
@@ 1.2 Significance and impact
Taken together, the framework supplies evaluation primitives for when and why global interpretability breaks, enabling atlas-level summaries that are reproducible, stability-tested, and comparable across subsystems, rather than anecdotal case studies.
@@ 11 Limitations
    Transport model class. We use linear transports on overlap codes; nonlinear transports may be required for certain strata.
@@ 12 Conclusion
We presented a discrete gauge-theoretic atlas formalism for superposition in LLMs and grounded it in four results with concrete empirical instantiations. Result A establishes a constructive gauge identity connecting chord residuals to holonomy, making holonomy computable and gauge-invariant. Result B turns proxy shearing into an unavoidable transfer mismatch bound. Result C provides a non-vacuous certified interference bound on consequential atom subsets with high coverage and zero violations across seeds/hyperparameters. Result D shows that the central diagnostics are stable under resampling and concentrate with overlap size, particularly on persistent subsystems. Together, these results support interpretability as an atlas construction problem with measurable local and global obstructions.

### v2

# A Gauge Theory of Superposition: Toward a Sheaf-Theoretic Atlas of Neural Representations

@@ Abstract
We prove and empirically instantiate four results on a frozen baseline open-weight LLM (Llama 3.2 3B Instruct) using standard text and code corpora (WikiText-103 raw, a C4-derived English web-text subset, and the-stack-smol) Meta (2024); Merity et al. (2016); Raffel et al. (2020); Kocetkov et al. (2022), and we further assess robustness through additional model-, layer-, seed-, null-, graph-, and transport-level ablations. (A) A constructive spanning-tree gauge theorem shows that, after gauge fixing on a spanning tree, each chord residual equals the holonomy of its fundamental cycle, making holonomy computable and gauge-invariant. (B) Shearing lower-bounds a data-dependent transfer mismatch energy, turning $D_{\mathrm{shear}}$ into an unavoidable failure bound. (C) A non-vacuous certified jamming/interference bound holds on consequential atom subsets with high coverage and zero violations across random seeds and hyperparameters. (D) Bootstrap and sample-size experiments establish estimator stability for $D_{\mathrm{shear}}$ and $D_{\mathrm{hol}}$, with improved concentration on persistent (well-conditioned) subsystems predicted by polar-factor perturbation theory Higham (1986); Schönemann (1966). Additional appendix experiments show that the diagnostics transfer across multiple model families and layers, remain stable across downstream seeds, respond strongly to a random-bases null control, and are robust to moderate changes in atlas, graph, and transport settings. Our framework supplies reproducible evaluation primitives for when and why global interpretability breaks, separating failures due to local overload (jamming), cross-chart mismatch (shearing), and intrinsically path-dependent transport (holonomy).
@@ 1.1 Contribution and paper thesis
To test whether these diagnostics reflect structural properties of learned representations rather than artefacts of one checkpoint or one implementation setting, the appendix reports a broader suite of robustness experiments. In particular, we report additional model-generalisation experiments (Qwen 2.5 3B Instruct Team (2024c; b) and Gemma 2 2B IT Team (2024a); Google (2024)), a layerwise sweep within the baseline Llama model, downstream seed-reproducibility tests, a random-bases null control, and robustness checks for atlas, graph, and transport hyperparameters. These experiments are intended to distinguish structural properties of the learned representation from artefacts of a single model or configuration.
@@ 1.2 Significance and impact
Taken together, the framework supplies evaluation primitives for when and why global interpretability breaks, enabling atlas-level summaries that are reproducible, stability-tested, and comparable across subsystems, rather than anecdotal case studies. The appendix further shows that these diagnostics are not confined to a single checkpoint or one hand-tuned configuration: they transfer across multiple model families, vary systematically across depth, remain stable across downstream seeds, respond strongly to a random-bases null control, and are robust to moderate changes in atlas, graph, and transport construction.
@@ 9.1 Experimental setup
The main text focuses on the core empirical instantiations of Results A–D on the baseline Llama 3.2 3B Instruct run. To assess whether these conclusions persist beyond a single checkpoint or a single implementation setting, the appendix reports additional robustness experiments covering cross-model transfer, layerwise variation, downstream seed reproducibility, a random-bases null control, atlas hyperparameter sweeps, graph-construction changes, and transport-regularisation ablations.
@@ 11 Limitations
    Transport model class. We use linear transports on overlap codes as a deliberately simple and reproducible baseline that keeps the transport estimation, polar-factor analysis, and holonomy computation analytically transparent. Nonlinear transports may be required for certain strata, and extending the framework to richer local map classes is an important direction for future work.
@@
  5. 5.
    Robustness scope. The appendix evaluates cross-model transfer, layerwise variation, downstream seed stability, a random-bases null control, and moderate atlas/graph/transport ablations. However, we do not exhaust all possible proxy choices, nonlinear transport classes, or downstream task-level operational validations, which remain directions for future work.
@@ 12 Conclusion
We presented a discrete gauge-theoretic atlas formalism for superposition in LLMs and grounded it in four results with concrete empirical instantiations. Result A establishes a constructive gauge identity connecting chord residuals to holonomy, making holonomy computable and gauge-invariant. Result B turns proxy shearing into an unavoidable transfer mismatch bound. Result C provides a non-vacuous certified interference bound on consequential atom subsets with high coverage and zero violations across seeds/hyperparameters. Result D shows that the central diagnostics are stable under resampling and concentrate with overlap size, particularly on persistent subsystems. Together, these results support interpretability as an atlas construction problem with measurable local and global obstructions. The appendix further shows that these diagnostics are not restricted to a single baseline run: they transfer across multiple model families, vary systematically across depth, remain stable across downstream seeds, respond strongly to a random-bases null control, and are robust to moderate changes in atlas, graph, and transport construction. This strengthens the case that the framework is not merely a formal reinterpretation, but an empirically executable methodology for diagnosing when and why global interpretability fails.